A chemist wants to make 700 mL of a 6% nitric acid solution by using 5% and 12% nitric acid solutions. How much of each should she use.
I got 699.4 mL of the 5% solution, and .06 mL of the 12% solution. It just doesn't seem right and I'd like to check.

Answer by TimothyLamb(4379)

---
5x + 12y = 6*700
x + y = 700
---
put the system of linear equations into standard form
---
5x + 12y = 4200
x + y = 700
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = 600 mL of 5%
y = 100 mL of 12%
---
system matrix
5 12
1 1
---
inverse of system matrix
-0.14285714 1.7142857
0.14285714 -0.71428571
---
determinant of system matrix = -7
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php